Kendra bikes to work. She travels at a rate of 12 miles per hour. How long will it take her to bike 8 miles
Answer by mananth(16946)

Time = distance / speed
distance == 8 miles
speed = 12 mph
time = 8/12 = 3/4 hour
=45 minutes